An automobile driver travels north for 2.0 min at 30 mph then he stops at a red light for 30 sec before proceeding again for 3.0 min at 45 mph. He then stops for a stop sign for 3.0 sec drives south at 30 mph for 2min and finally stops for gas. What is the average velocity in mile and minute when the positive x direction is north
Answer #1

s1 = u1 x t1 = 30 x (2/60) = 1 mi
=>s2 = u2 x t2 = 0 x (30/3600) = 0 mi
=>s3 = u3 x t3 = 45 x (3/60) = 2.25 mi
=>s4 = u4 x t4 = 0 x (3/3600) = 0 mi
=>s5 = u5 x t5 = 30 x (2/60) = 1 mi
v(avg) = [s1 + s2 + s3 + s4 + s5] / [t1+t2+t3+t4+t5]
=>v(avg) = [1 + 0 + 2.25 + 0 + 1] / [(2/60) + (30/3600) + (3/60) + (3/3600) + (2/60)]
=>v(avg) = 4.25/0.126 = 33.8 mi/hr